OpenStudy (anonymous):

Write log315 - log310 as a single logarithm A. log33 B. log3 3/2 C. log32 D. log33 - log32

OpenStudy (anonymous):

\[\log_{3} 15 - \log_{3}10\ = log_{3} \frac{15}{10} = log_{3} \frac{3}{2}\]Using the law of logarithms loga - logb = log(a/b).